OpenStudy (anonymous):

NEED HELP!! Solve using the square root property. 3a^2-20=13

OpenStudy (anonymous):

\[\large 3a^2-20=13 \rightarrow 3a^2=27 \rightarrow a^2=9\] can you solve from here?

OpenStudy (anonymous):

\[3a^2-20=13\]\[3a^2=13+20\]\[3a^2=33\]

OpenStudy (anonymous):

I got 3a^2=33, is there anything more or is that the final answer?

OpenStudy (anonymous):

nope \[a^2 =33/11 \] \[a=\pm \sqrt{33/11}\]

OpenStudy (anonymous):

=1.732

OpenStudy (anonymous):

ok, thank you :o)

OpenStudy (anonymous):

no problem :D